Tohban Report 2014-11-05
From RHESSI Wiki
|Start Date:||29 Oct 2014|
|End Date:||05 Nov 2014|
|Next Tohban:||Juan Carlos Martinez-Oliveros|
|List all reports|
Solar activity began high with the gigantic AR 12192, but then tapered off as the region rotated over the limb. Activity has picked up in the past 48 hours as AR 12205 rotated on disk and produced a series of large M-class flares. This spot is expected to produce some additional large events in the coming days. With the exception of the series of X-class events, the general flaring activity this week was comparable to last week.
How many GOES flares occurred?
Flares above B, C, M, X class were 3 34 14 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 0 22 13 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 11 5 0
There were RHESSI flares/GOES flares 324 / 51 over the time range 22-Oct-14 29-Oct-14
SSR emptied by mid-week and has been continuing to empty at the end of the 2nd to last pass. We remain in Active/Vigorous mode and night time time data collection remained off throughout the week. This coming week we will turn on full rear eclipse events.
All temperatures are nominal this week, cold plate/tip 1 are at 125/104.9K and cold plate/tip 2 are at 123/103K.
Attenuator motion (80 thin and 2 thick motions) and cryocooler was nominal. We hit 70,000 orbits over the weekend.
All gaps were replayed and should be filled in
GAPS IN APP_ID = 1 (VC1-SOH)
2014-11-03T06:45:00.000 -- 2014-11-03T06:50:00.000 300.00000
2014-11-03T07:35:00.000 -- 2014-11-03T07:40:00.000 300.00000
2014-11-03T08:30:00.000 -- 2014-11-03T08:45:00.000 900.00000
2014-11-03T15:35:00.000 -- 2014-11-03T15:50:00.000 900.00000
GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4
2014-11-03T15:35:00.000 -- 2014-11-03T15:45:00.000 600.00000
GAPS IN APP_ID = 102 (VC3-MONITOR RATES)
2014-11-02T02:20:00.000 -- 2014-11-02T04:10:00.000 6600.0000
Detector 2: Light leak issues persist
Detector 9: Over the time period of Oct 31 - Nov 2, D9's monitor rates plots show elevated noise. The onset was sudden, as was the end. There was a 45 minute period on Nov 1st, from 19:00 - 19:45, when the noise returned to normal values. Since Nov 2nd, D9 has had normal levels of noise. There are currently no theories as to why this occurred.
There were no threshold changes this week.
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||continue +/- 4 minutes|
|Night time data (rears)||turn-on Full nighttime events|
|Require extra passes?||Yes|
|Requirement for moving pointer?||No|
|Detector problems?||Detector 2 light leak. Detector 9 elevated noise Oct 31 - Nov 2.|